If you inoculate 1*10^6 CFU/ml E.coli, assuming a population doubling time of 20 minutes. Calculate the total number of cells present in a 100 ml flask of LB broth after 6 hours.
1
Expert's answer
2019-08-14T05:56:51-0400

Initially, the number of cells was 1*10^6  CFU/ml * 100 ml = 1*10^8  CFU

The number of 20-minutes intervals in 6 hours is 18.

The expected number of bacteria is 2^18*1*10^8 = 262144*10^8. It is approximately 2.6*10^13 cells